OpenStudy (raden):

Find all integer number (positive) for n, such that : n^4 - 4n^3 + 22n^2 - 36n + 18 are perfect square numbers!

OpenStudy (raden):

@mukushla, help me

OpenStudy (anonymous):

i will if i can

OpenStudy (raden):

i got one of solution is n=1, but any else i dont know....

OpenStudy (anonymous):

man wait ... im getting something interesting

OpenStudy (anonymous):

\[n^4 - 4n^3 + 22n^2 - 36n + 18=(n^2-2n+9)^2-63\]

OpenStudy (anonymous):

\[(n^2−2n+9)^2−63=a^2-63=b^2\]\[a^2-b^2=63\]it is possible only when \(a=8\) and \(b=1\)

OpenStudy (anonymous):

so \(n=1\) is the only answer

OpenStudy (raden):

hmmm, very interesting ur way for me... because i got my answer with trial and error ^^ but after i see your answer, mukushla i got the answer else : a^2 - b^2 = 63 (with a=n^2-2n+9) (a+b)(a-b) = 63 i take 2 case : => for the case : (a+b)(a-b) = 9*7 so, a+b=9 and a-b=7 (or reverse) i got a=8 (got n=1) => for the case : (a+b)(a-b) = 3*21 so, a+b=3 and a-b=21 (or reverse) i got a=12 thus, n^2-2n+9 = 12 n^2-2n-3 = 0 (n-3)(n+1) = 0 n=3 (n=-1 cant) So, the solutions for n is {1,3) Isn't right mukushla?

OpenStudy (anonymous):

oh yes thats right...silly mistake from my side

OpenStudy (raden):

alhamdulillah, thanks for ur help mukushla... i cant solve this problem, if i dont see ur step the first... hehe

OpenStudy (anonymous):

np man...:)